# Abmax: A JAX-based Agent-based Modeling Framework

### v1

@@
Abstract
Agent-based modeling (ABM) is a principal approach for studying complex systems. By decomposing a system into simpler, interacting agents, agent-based modeling (ABM) allows researchers to observe the emergence of complex phenomena. High-performance array computing libraries like JAX can help scale such computational models to a large number of agents by using automatic vectorization and just-in-time (JIT) compilation. One of the caveats of using JAX to achieve such scaling is that the shapes of arrays used in the computational model should remain immutable throughout the simulation. In the context of agent-based modeling (ABM), this can pose constraints on certain agent manipulation operations that require flexible data structures. A subset of which is represented by the ability to update a dynamically selected number of agents by applying distinct changes to them during a simulation. To this effect, we introduce Abmax, an ABM framework based on JAX that implements multiple just-in-time (JIT) compilable algorithms to provide this functionality. On the canonical predation model benchmark, Abmax achieves runtime performance comparable to state-of-the-art implementations. Further, we show that this functionality can also be vectorized, making it possible to run many similar agent-based models in parallel. We also present two examples in the form of a traffic-flow model and a financial market model to show the use case of Abmax. [^1] Abmax is available at https://github.com/i-m-iron-man/abmax.git
Keywords: Agent-based Models, JAX, Complex Systems
1 Introduction
Complex systems are systems that are composed of many interacting parts. Examples of such systems include atoms in a molecule (Testa and Bojarski, 2000), neurons in a brain (Kulkarni and Bassett, 2025), companies in a market (Foster, 2005), computers on the internet (Park, 2005), or people in a population (Rutter et al., 2017). It is important to model, simulate, and analyze such systems to uncover the properties of complex phenomena that can emerge in them (Kwapień and Drożdż, 2012). Examples are phase changes in matter, emergence of intelligence in brains, crashing of financial markets, spread of a computer virus through the internet, or spread of an infectious disease within a population.
@@ 1 Introduction
With the increased availability of compute power in the form of hardware accelerators such as GPUs and TPUs, it has become possible to simulate computational models of a large number of agents (Xiao et al., 2019). This helps to reduce the gap between these models and the actual phenomena they are trying to model. To facilitate agent-based modeling and simulation, various software frameworks have been developed. Notable examples are the Python-based Mesa framework (Kazil et al., 2020), the Julia-based Agents.jl framework (Datseris et al., 2022) and the C++-based Flame-GPU framework (Richmond and Chimeh, 2017).
In this work, we introduce Abmax as a novel, generic Python-based ABM framework, combining ease of use with computational efficiency. The former is provided by an intuitive Python application programming interface (API), whereas the latter is provided by using JAX Bradbury et al. (2018) as the backend engine. JAX is a Python native high-performance array computing library that provides automatic vectorization across accelerators and just-in-time (JIT) compilation through simple decorators, allowing modellers to develop efficient large-scale models entirely in Python. This has been extensively showcased in the field of reinforcement learning (Lu et al., 2022) and computational simulations of evolution (Lange, 2023). Abmax provides the research community with these capabilities to support large-scale agent-based modeling.
The structure of this paper is as follows. Section 2 introduces the Abmax API. Section 3 describes implementational details of how Abmax achieves efficient simulation of many-agent models with a dynamically varying number of agents. In Section 4, we compare Abmax to other agent-based modeling frameworks and demonstrate scaling towards large-scale simulations. Finally, in the Discussion section, we present limitations and future work.
2 Abmax
@@
Finally, a Set is a wrapper around a collection of Agents. It stores an array of Agent objects and two integers num_agents and num_active_agents that represent the maximum number of agents possible and the current number of agents in the set, respectively. Further, it also mirrors most of the fields held by a single Agent, including a State, Params, and Policy object. This structure makes it easier to run batched and cascaded ABM simulations using a single vmap or pmap.
Figure 1: Unified Modeling Language (UML) diagram representing different classes in the Abmax framework. These classes can be divided into the helper and the core classes.
@@
3 Implementation
In many agent-based modeling settings, being able to apply distinct updates to a dynamically chosen, variable-size subset of agents is a crucial feature needed for accurate ABM. This includes cases where the number of agents in the simulations is dynamic and new agents need to be added to the simulation with distinct initial states and parameters.
@@ 3 Implementation
Abmax provides the ability to add, remove, or update a dynamically selected number of agents that can be determined during runtime through a JIT compiled function. This capability can further be completely vectorized across homogeneous data-structures, allowing the parallel simulation of multiple models. The main constraint of our approach is that modelers must know the maximum number of agents in advance, since we use placeholder agents to keep data structures’ shapes constant throughout the simulation. Within the Abmax framework, we present Rank-Match and Sort-Count-Iterate as two algorithms to realize this functionality, with the former focusing on speed and the latter focusing on modeling flexibility.
@@ 4.1 Predation model
Figure 2: Depiction of Wilensky’s discrete 2D predation model (Wilensky, 1997). Wolves gain energy by consuming sheep, and sheep gain energy by eating grass, which regrows on each cell after a fixed delay. Both species move randomly and reproduce asexually with a given probability, spawning offspring at their current location. Abmax leverages JAX’s vmap to run multiple such simulations in parallel over identical data structures.
@@
Table 1: Time taken in simulating 100 steps of the predation model reported in milliseconds for small and large environments. The small environment starts with 600 sheep and 400 wolves spawned on a 100 by 100 lattice. A large environment starts with 6000 sheep and 4000 wolves on a 1000 by 1000 lattice. The times are calculated by subtracting the median runtime of five steps from that of 105 steps over ten runs to erase the effect of initial setup time, which is substantially high for JAX.
Environment size Agents.jl Abmax RM Abmax SCI Mesa
small 14.93 50.26 726.78 1333.047
large 685.03 3315.88 5455.01 170070.95
Another useful feature of Abmax is that it can run multiple agent-based models in parallel over hardware accelerators by simply using vmap or pmap from JAX. The speed-up enabled by this feature is visible in Table 2.
Table 2: Time taken in simulating 100 steps of multiple predation models reported in milliseconds. The parameters of this model are the same as the small environment in Table 1.
number of models 10 20 50 100 200 500
time taken (ms) 3088.58 3171.38 3083.56 3316.24 3616.16 5343.62
An example of such a simulation can be seen in Figure 3, where ten populations of wolves and sheep were simulated in ten different environments in parallel. These simulations start with similar initial conditions but different seeds, and thus the difference in trend reflects the stochasticity in the position and reproduction of agents.
Figure 3: Trend in the number of wolves and number of sheep for similar initial conditions using ten different seeds, simulated for 100 steps in a small predation environment.
4.2 Traffic simulation
@@
Figure 4: A cellular‐automaton traffic model: green arrows mark allowed moves, red arrows blocked actions, and purple arrows potential conflicts. Cars spawn in yellow entry cells and travel toward exit cells under green/red signals. By representing everything with Abmax data structures and applying JAX’s vmap, we can parallelize simulations across multiple roads with different car counts.
@@
Table 3: Time taken in simulating 1000 steps of the traffic-flow simulation for different numbers of roads. The roads are three lanes wide and 100 cells long, thus capable of housing 300 cars at a time. It can be seen that the vectorization across the roads helps prevent the runtime from scaling in proportion to the number of roads.
number of roads 10 20 50 100 200 500
time taken (ms) 931.65 996.60 1028.15 1230.85 1467.89 2327.71
@@
Figure 5: The variation in the number of cars across ten different roads for 100 steps. Wherein each road has a different seed leading to different trends in the green/red signals in the exit cells, and a different number of cars are spawned at each step. Each road is seven cells long and three lanes wide, thus it can host a maximum of 21 cars.
4.3 Financial market
@@
Figure 6: Noisy traders place buy or sell orders across multiple limit order books (LOBs), where orders may be active, inactive, or new. After each step, when all traders have placed their orders, all LOBs are matched in parallel. Using Abmax data structures and JAX’s vmap, this matching is vectorized across every book.
@@
Table 4: Time taken in simulating 100 steps of finance markets with different numbers of limit order books (LOBs) of maximum size 1000 each.
number of LOBs 10 20 50 100 200 500
time taken (ms) 297.01 333.47 470.88 663.80 1034.36 2212.47
@@
Figure 7: The (a) market price, (b) number of active buy orders, and (c) number of active sell orders across five limit order books (LOBs) are simulated for 100 steps. The maximum size of each LOB is 1000.
5 Discussion
@@
Using Abmax, we also developed algorithms that are specific to the context of traffic modeling using cellular-automata and limit order book manipulation in a financial market. While developing them, one common theme that kept reappearing was that brute-force-based approaches, when vectorized, often outperformed an approach optimized for serial computations. This observation brings to light the fact that, with the rise in popularity of hardware accelerators, we might need to revisit how we formulate algorithms. The goal in such an exercise would be to investigate if a vectorized brute-force variant of the formulation exists and if it can, by virtue of vectorization, outperform the streamlined variant designed for serial computations. Through Abmax, we aim at maintaining a collection of such algorithms focused on ABM that are also friendly to JAX transforms.

### v3

@@
Abstract.
Agent-based modeling (ABM) is a principal approach for studying complex systems. By decomposing a system into simpler, interacting agents, agent-based modeling (ABM) allows researchers to observe the emergence of complex phenomena. High-performance array computing libraries like JAX can help scale such computational models to a large number of agents by using automatic vectorization and just-in-time (JIT) compilation. One of the caveats of using JAX to achieve such scaling is that the shapes of arrays used in the computational model should remain immutable throughout the simulation. In the context of agent-based modeling (ABM), this can pose constraints on certain agent manipulation operations that require flexible data structures. A subset of which is represented by the ability to update a dynamically selected number of agents by applying distinct changes to them during a simulation. To this effect, we introduce ABMax, an ABM framework based on JAX that implements multiple just-in-time (JIT) compilable algorithms to provide this functionality. On the canonical predation model benchmark, ABMax achieves runtime performance comparable to state-of-the-art implementations. Further, we show that this functionality can also be vectorized, making it possible to run many similar agent-based models in parallel. We also present two examples in the form of a traffic-flow model and a financial market model to show the use case of ABMax.
Key words and phrases:
Agent-based Models, JAX, Complex systems
1. Introduction
Complex systems are systems that are composed of many interacting parts. Examples of such systems include atoms in a molecule (Testa and Bojarski, 2000), neurons in a brain (Kulkarni and Bassett, 2025), companies in a market (Foster, 2005), computers on the internet (Park, 2005), or people in a population (Rutter et al., 2017). It is important to model, simulate, and analyze such systems to uncover the properties of complex phenomena that can emerge in them (Kwapień and Drożdż, 2012). Examples are phase changes in matter, the emergence of intelligence in brains, the crashing of financial markets, the spread of a computer virus through the internet, or the spread of an infectious disease within a population.
@@ 1. Introduction
With the increased availability of compute power in the form of hardware accelerators such as Graphics Processing Units (GPU) and Tensor Processing Units (TPU), it has become possible to simulate computational models of a large number of agents (Xiao et al., 2019). This helps to reduce the gap between these models and the actual phenomena they are trying to model. To facilitate agent-based modeling (ABM) and simulation, various software frameworks have been developed. Notable examples are the Python-based Mesa framework (Kazil et al., 2020), the Julia-based Agents.jl framework (Datseris et al., 2022) and the C++-based Flame-GPU framework (Richmond and Chimeh, 2017).
In many agent-based modeling settings, being able to apply distinct updates to a dynamically chosen, variable-size subset of agents is a crucial feature for accurate ABM. This includes cases where the number of agents in the simulation is dynamic and new agents must be added with distinct initial states and parameters, or where multiple agents opt for the same resource, requiring a conflict-resolution protocol (Yang et al., 2018). In this work, we introduce ABMax, a Python-based ABM framework built on JAX (Bradbury et al., 2018) that provides two algorithms for applying such updates under JAX’s static-shape constraints. This capability can further be completely vectorized across homogeneous data structures, allowing the parallel simulation of multiple agent-based models.
ABMax combines ease of use with computational efficiency. The former is provided by an intuitive Python application programming interface (API), whereas the latter is provided by using JAX as the backend engine. JAX is a Python native high-performance array computing library that provides automatic vectorization across accelerators and just-in-time (JIT) compilation through simple decorators, allowing modelers to develop efficient large-scale models entirely in Python. This has been extensively showcased in the field of reinforcement learning (Lu et al., 2022) and computational simulations of evolution (Lange, 2023). ABMax provides the research community with these capabilities to support large-scale agent-based modeling.
The structure of this paper is as follows. Section 2 introduces the ABMax API. Next, in Section 3 we describe the general design structure of an agent-based model developed using ABMax. Section 4 describes implementational details of two algorithms, namely Rank-Match (RM) and Sort-Count-Iterate (SCI) used within ABMax to provide distinct updates to a subset of agents selected at runtime. In Section 5, we compare ABMax to other agent-based modeling frameworks and demonstrate scaling towards large-scale simulations. Finally, in the Discussion section, we present limitations and future work.
2. ABMax
@@
Finally, a Set is a wrapper around a collection of Agents. It stores an array of Agent objects and two integers num_agents and num_active_agents that represent the maximum number of agents possible and the current number of agents in the set, respectively. Further, it also mirrors most of the fields held by a single Agent, including a State, Params, and Policy object. This structure makes it easier to run batched and cascaded ABM simulations using a single vmap.
Figure 1. Unified Modeling Language (UML) diagram representing different classes in the ABMax framework. These classes can be divided into the helper and the core classes.
@@
3. Simulation Design
A typical design of a project based on ABMax can be seen below.
import jax
import abmax
from flax import struct
class Agent_1(abmax.structs.Agent):
def create_agent(…) -> abmax.structs.Agent:
def step_agent(…) -> abmax.structs.Agent:
def update_agent(…) -> abmax.structs.Agent:
class Agent_2(abmax.structs.Agent):
@jax.jit
def agent_interactions(…):
Class Sim():
Agent_A1_set: abmax.structs.Set
Agent_A2_set: abmax.structs.Set
def create_sim(…) -> Sim:
def step(…) -> Sim:
def main(…):
sim = Sim.create_sim(…)
sim = jax.lax.scan(Sim.step)(sim, …)
Each heterogeneous agent class can be defined by inheriting the Agent class from ABMax. The class encapsulates a create_agent function that contains the procedure to initialize a single agent; this function is later passed to create_agents for instantiating a vectorized data structure of agents. Similarly, the step_agent member implements the logic to step a single agent in the simulation, which is vectorized across all agents of that class by a JIT-compiled callback to step_agents. The update_agent member defines single-agent update logic that is applicable to a subset of agents. Depending on whether the applicability is unique or common to all agents in the subset, this member is passed as an argument to either the set_agents_rm and set_agents_sci kernels (for unique updates) or to the set_agents_mask kernel (for non-unique updates).
Next, the agent_interactions function represents the functions that define how different agents interact with each other. Such functions differ across ABM contexts and thus lend themselves to a set of design patterns. With ABMax, we aim to maintain a collection of these patterns. A separate class Sim can be declared to set up all the agent sets in the project and to define how the entire simulation advances one step via the members create_sim and step, respectively. Agent-set manipulation functions like set_agents_sci, set_agents_rm, set_agents_mask, select_agents, and
sort_agents can be used in Sim.step or agent_interactions to achieve the desired agent manipulation.
In practice, all classes are decorated with the
flax.struct.dataclass decorator so they comply with JAX transforms, and relevant class member functions are written as
staticmethods to align with the functional programming paradigm used with JAX. Finally, Sim objects can be scanned using jax.lax.scan for the desired number of time steps to run the simulation.
4. Implementation
ABMax provides the ability to update a dynamically selected number of agents that can be determined during runtime through a JIT compiled function. This capability can also be completely vectorized across homogeneous data structures, allowing the parallel simulation of multiple models wherein a different number of agents can be updated distinctly in each of them. The main constraint of our approach is that modelers must know the maximum number of agents in advance, since we use placeholder agents to keep data structures’ shapes constant throughout the simulation. Within the ABMax framework, we present Rank-Match and Sort-Count-Iterate as two algorithms to realize this functionality, with the former focusing on speed and the latter focusing on modeling flexibility.
@@ 5.1. Predation model
Figure 2. Depiction of Wilensky’s discrete 2D predation model (Wilensky, 1997). Wolves gain energy by consuming sheep, and sheep gain energy by eating grass, which regrows on each cell after a fixed delay. Both species move randomly and reproduce asexually with a given probability, spawning offspring at their current location. ABMax leverages JAX’s vmap to run multiple such simulations in parallel over identical data structures.
@@
Table 1. Time taken in simulating 100 steps of the predation model reported in milliseconds for small and large environments. The small environment starts with 600 sheep and 400 wolves spawned on a 100 by 100 lattice. A large environment starts with 6000 sheep and 4000 wolves on a 1000 by 1000 lattice. The times are calculated by subtracting the median runtime of five steps from that of 105 steps over ten runs to erase the effect of initial setup time, which is substantially high for JAX.
Environment size small large
Agents.jl $14.93$ $685.03$
ABMax RM $50.26$ $3315.88$
ABMax SCI $726.78$ $5455.01$
Mesa $1333.047$ $170070.95$
Another useful feature of ABMax is that it can run multiple agent-based models in parallel over hardware accelerators by simply using vmap from JAX. The speed-up enabled by this feature is visible in Table 2.
Table 2. Time taken in simulating 100 steps of multiple predation models reported in milliseconds. The parameters of this model are the same as the small environment in Table 1.
number of models time taken (ms)
$10$ $3088.58$
2020 $3171.38$
$50$ $3083.56$
$100$ $3316.24$
$200$ $3616.16$
$500$ $5343.62$
An example of such a simulation can be seen in Figure 3, where ten populations of wolves and sheep were simulated in ten different environments, all in parallel. These simulations start with similar initial conditions but with different seeds, and thus the difference in trend reflects the stochasticity in the position and reproduction of agents.
Figure 3. Trend in the number of wolves and number of sheep for similar initial conditions using ten different seeds, simulated for 100 steps in a small predation environment.
5.2. Traffic simulation
@@
Figure 4. A cellular‐automaton traffic model: green arrows mark allowed moves, red arrows blocked actions, and purple arrows potential conflicts. Cars spawn in yellow entry cells and travel toward exit cells under green/red signals. By representing everything with ABMax data structures and applying JAX’s vmap, we can parallelize simulations across multiple roads with different car counts.
@@
Table 3. Time taken in simulating 1000 steps of the traffic-flow simulation for different numbers of roads. The roads are three lanes wide and 100 cells long, thus capable of housing 300 cars at a time.
number of roads time taken (ms)
$10$ $931.65$
2020 $996.60$
$50$ $1028.15$
$100$ $1230.85$
$200$ $1467.89$
$500$ $2327.71$
@@
Figure 5. The variation in the number of cars across ten different roads for 100 steps. Wherein each road has a different seed leading to different trends in the green/red signals in the exit cells, and a different number of cars are spawned at each step. Each road is seven cells long and three lanes wide, thus it can host a maximum of 21 cars.
5.3. Financial market
@@
Figure 6. Noisy traders place buy or sell orders across multiple limit order books (LOBs), where orders may be active, inactive, or new. After each step, when all traders have placed their orders, all LOBs are matched in parallel. Using ABMax data structures and JAX’s vmap, this matching is vectorized across every book.
@@
Table 4. Time taken in simulating 100 steps of finance markets with different numbers of limit order books (LOBs) of maximum size 1000 each.
number of LOBs time taken (ms)
$10$ $297.01$
2020 $333.47$
$50$ $470.88$
$100$ $663.80$
$200$ $1034.36$
$500$ $2212.47$
@@
Figure 7. The (a) market price, (b) number of active buy orders, and (c) number of active sell orders across five limit order books (LOBs) are simulated for 100 steps. The maximum size of each LOB is 1000.
6. Discussion
@@
Using ABMax, we also developed algorithms that are specific to the context of traffic modeling using cellular-automata and limit order book manipulation in a financial market. While developing them, one common theme that kept reappearing was that brute-force-based approaches, when vectorized, often outperformed an approach optimized for serial computations. This observation brings to light the fact that, with the rise in popularity of hardware accelerators, we might need to revisit how we formulate algorithms. A similar observation was recently noted in the field of genetic-programming (De Vries et al., 2025). The goal in such an exercise would be to investigate if a vectorized brute-force variant of the formulation exists and if it can, by virtue of vectorization, outperform the streamlined variant designed for serial computations. Through ABMax, we aim at maintaining a collection of such algorithms focused on ABM that are also friendly to JAX transforms.